OpenStudy (anonymous):

Write the logarithm in expanded form. ln (r to 9s over t to the 6

OpenStudy (anonymous):

\[\ln (r^9s/t^6)\]

OpenStudy (anonymous):

9ln(r) + ln(s) - 6ln(t)

OpenStudy (anonymous):

\[\log base of 9 (x/y^6z^4)\]

OpenStudy (anonymous):

Here, when I write log, I mean log base of 9. The answer: log x - 6log y -4log z

OpenStudy (anonymous):

how come it is all subtract?

OpenStudy (anonymous):

Distributive property.

OpenStudy (anonymous):

\[\log \to the base 7 (4 on the \top thing \sqrt{xz}\]

OpenStudy (anonymous):

Clarify, please?

OpenStudy (anonymous):

 Write the logarithm in expanded form.